### v3

# Anton Alexa

Abstract
We establish a boundary delocalization principle for high–frequency Dirichlet eigenfunctions on smooth strictly convex domains. The main result is an exclusion of persistent boundary concentration at the level of individual eigenmodes relative to short spectral packets of sublinear length. Quantitatively, we compare boundary energies of single eigenfunctions to packet sums over frequency windows of size $N_{k}=o(k)$, without asserting any asymptotic gain in magnitude. The main mode–to–packet estimate relies only on the Rellich identity. For the multi–mode bias exclusion we additionally use the boundary local Weyl law to obtain a packet zero–mean cancellation estimate. This mode–to–packet comparison is independent of eigenvalue monotonicity and is stable under eigenvalue crossings.
1 Introduction
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded $C^{\infty}$ strictly convex domain, $d\geq 2$, and let $\{(\lambda_{k},u_{k})\}_{k\geq 1}$ denote the Dirichlet eigenpairs of $\Delta_{\Omega}$ with $\|u_{k}\|_{L^{2}(\Omega)}=1$. The boundary flux density $|\partial_{n}u_{k}|^{2}$ plays a central role in boundary spectral analysis and in microlocal restriction phenomena. A natural question is whether a single high–frequency eigenfunction can exhibit dominant boundary localization compared to its spectral neighbors.
Related results on boundary delocalization and restriction estimates typically rely on microlocal analysis, defect measures, or dynamical properties of the billiard flow [5, 6, 12, 7, 1]. For general background on boundary eigenfunctions and related phenomena [14, 15]. In contrast, the main mode–to–packet estimate is entirely static and relies only on the Rellich identity. For the multi–mode bias exclusion we additionally use the boundary local Weyl law to obtain packet zero–mean cancellation.
This paper shows that, on strictly convex domains, persistent boundary concentration at the level of a single mode is excluded when one compares it to a spectral packet of sublinear length. We establish a packet–level delocalization principle: for any sublinear window $N_{k}=o(k)$, the ratio between a single boundary energy and the packet sum tends to zero. The proof uses the Rellich identity. In addition, we rule out collective boundary concentration: no positive fraction of modes inside a sublinear spectral packet can exhibit a persistent one–sided boundary bias with a fixed geometric weight of zero mean, in the spirit of averaged spectral quantities; [11, 10]. For the multi–mode statement we assume a slightly stronger growth condition on $N_{k}$ when $d\geq 3$.
Theorem 1.1 (Main result - Introductory Form).
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded $C^{\infty}$ strictly convex domain, $d\geq 2$, and let $\{(\lambda_{k},u_{k})\}_{k\geq 1}$ be the Dirichlet eigenpairs of $\Delta_{\Omega}$. Let $N_{k}$ be any sublinear spectral window with $N_{k}\to\infty$ and $N_{k}=o(k)$. Then
$\frac{E_{k}}{\sum_{m=k}^{k+N_{k}-1}E_{m}}\longrightarrow 0\quad\text{as }k\to\infty.$ (1)
In particular, no individual high–frequency eigenfunction can asymptotically carry a fixed positive proportion of the boundary energy of a sublinear spectral packet.
Remark 1.2 (Sharpness of the packet length condition).
The condition $N_{k}\to\infty$ is essential at the level of the present method. Indeed, for a fixed packet length $N$, the Rellich identity and the Weyl law alone do not imply that
$\frac{E_{k}}{\sum_{m=k}^{k+N-1}E_{m}}\longrightarrow 0.$ (2)
In particular, without averaging over a growing spectral window, a single eigenmode may contribute a non–negligible fraction of the packet boundary energy. Thus the requirement $N_{k}\to\infty$ represents the minimal averaging scale necessary to enforce mode–to–packet delocalization.
2 Preliminaries and notation
Let $\Omega\subset\mathbb{R}^{d}$ be bounded, $C^{\infty}$, and strictly convex. Denote by $\{(\lambda_{k},u_{k})\}_{k\geq 1}$ the Dirichlet eigenpairs of $\Delta_{\Omega}$, with $u_{k}|_{\partial\Omega}=0$ and $\|u_{k}\|_{L^{2}(\Omega)}=1$. For $x\in\partial\Omega$, define the boundary flux density
$\rho_{k}(x):=|\partial_{n}u_{k}(x)|^{2}.$ (3)
The total boundary energy is
$E_{k}:=\int_{\partial\Omega}\rho_{k}(x)\,d\sigma(x).$ (4)
Let $H$ denote the mean curvature of $\partial\Omega$ with respect to the outer normal and let $\bar{H}$ be its boundary average.
Definition 2.1 (Spectral packets).
Let $N_{k}$ be a positive integer sequence with $N_{k}\to\infty$ and $N_{k}=o(k)$. Define the index packet
$\mathcal{J}_{k}:=\{k,\,k+1,\,\dots,\,k+N_{k}-1\}$ (5)
and the packet boundary density
$\mathcal{Q}_{k,N_{k}}(x):=\frac{1}{N_{k}}\sum_{j\in\mathcal{J}_{k}}\rho_{j}(x).$ (6)
For a measurable weight $w:\partial\Omega\to\mathbb{R}$, set
$E_{k}(w)$ $=\int_{\partial\Omega}\rho_{k}(x)\,w(x)\,d\sigma(x),$ (7)
$E_{k}^{abs}(w)$ $=\int_{\partial\Omega}\rho_{k}(x)\,|w(x)|\,d\sigma(x),$
$\mathcal{E}_{k,N_{k}}(w)$ $=\frac{1}{N_{k}}\sum_{j\in\mathcal{J}_{k}}E_{j}^{abs}(w).$
Definition 2.2 (Boundary correlation coefficient).
Let $w\in L^{\infty}(\partial\Omega)$ satisfy $\int_{\partial\Omega}w\,d\sigma=0$. Define
$C_{k}(w):=\frac{1}{E_{k}}\int_{\partial\Omega}\rho_{k}(x)\,w(x)\,d\sigma(x).$ (8)
For the multi–mode statement we assume $w\in C^{\infty}(\partial\Omega)$.
Remark 2.3 (Eigenvalue crossings).
All arguments are formulated at the level of boundary energy densities $\rho_{k}=|\partial_{n}u_{k}|^{2}$ and spectral sums. No ordering or differentiability of individual eigenvalues is required, and possible eigenvalue crossings do not affect the packet–level estimates.
3 Rellich identity and single–mode bounds
The Rellich identity yields the correct scaling for the boundary energy of a single eigenfunction on strictly convex domains; [13, 3].
Lemma 3.1 (Single–mode boundary energy bound).
There exist constants $0<c<C$ such that for all $k\geq 1$,
$c\,\lambda_{k}\leq E_{k}\leq C\,\lambda_{k}.$ (9)
Moreover, for any $w\in L^{\infty}(\partial\Omega)$,
$|E_{k}(w)|\leq E_{k}^{abs}(w)\leq C\,\lambda_{k}\,\|w\|_{L^{\infty}(\partial\Omega)}.$ (10)
Proof 3.2.
Fix $x_{0}\in\Omega$ and set $g(x):=\langle x-x_{0},\nu(x)\rangle$ on $\partial\Omega$. Since $\Omega$ is convex, for each $x\in\partial\Omega$ one has $\langle y-x,\nu(x)\rangle\leq 0$ for all $y\in\Omega$ (supporting hyperplane), where $\nu(x)$ is the outward unit normal. Since $x_{0}$ is interior, it cannot lie on the supporting hyperplane at $x$, hence $\langle x_{0}-x,\nu(x)\rangle<0$, and thus $g(x)=\langle x-x_{0},\nu(x)\rangle>0$ for all $x\in\partial\Omega$. By continuity and compactness, $0<m\leq g\leq M<\infty$ on $\partial\Omega$. The Rellich identity [13] yields
$\int_{\partial\Omega}g(x)\,|\partial_{n}u_{k}(x)|^{2}\,d\sigma(x)=2\lambda_{k}.$ (11)
Therefore
$\frac{2}{M}\lambda_{k}\leq E_{k}\leq\frac{2}{m}\lambda_{k},$ (12)
which gives the claimed bounds. The weighted estimate follows from $|E_{k}(w)|\leq E_{k}^{abs}(w)\leq\|w\|_{L^{\infty}}E_{k}$. ∎
4 Packet scales and a contextual Weyl remark
The integrated boundary energy admits a Weyl-type leading asymptotic on smooth strictly convex domains; [2, 4] We record this only for context.
Definition 4.1 (Integrated boundary energy).
Define the partial boundary energy sum
$S(\Lambda):=\sum_{\lambda_{j}<\Lambda}E_{j}.$ (13)
Remark 4.2 (Integrated boundary Weyl expansion (context)).
It is known that
$S(\Lambda)=C_{\Omega}\,\Lambda^{1+\frac{d}{2}}+o\!\left(\Lambda^{1+\frac{d}{2}}\right),$ (14)
as $\Lambda\to\infty$; [2, 4, 9]. This is recorded for context and is not used in the proof of the main theorem.
Lemma 4.3 (Packet zero–mean cancellation).
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded $C^{\infty}$ strictly convex domain. Let $w\in C^{\infty}(\partial\Omega)$ satisfy $\int_{\partial\Omega}w\,d\sigma=0$. Let $N_{k}\to\infty$ with $N_{k}=o(k)$ and $N_{k}/k^{1-2/d}\to\infty$, and set $\mathcal{J}_{k}=\{k,\dots,k+N_{k}-1\}$. Then
$\sum_{j\in\mathcal{J}_{k}}\int_{\partial\Omega}|\partial_{n}u_{j}(x)|^{2}\,w(x)\,d\sigma(x)=o\!\left(\sum_{j\in\mathcal{J}_{k}}E_{j}\right)\qquad(k\to\infty).$ (15)
Equivalently,
$\int_{\partial\Omega}\mathcal{Q}_{k,N_{k}}(x)\,w(x)\,d\sigma(x)=o\!\left(\frac{1}{N_{k}}\sum_{j\in\mathcal{J}_{k}}E_{j}\right)\qquad(k\to\infty).$ (16)
Proof 4.4.
For $\Lambda>0$ define the boundary spectral density
$Q_{\Lambda}(x):=\sum_{\lambda_{j}<\Lambda}|\partial_{n}u_{j}(x)|^{2},\qquad x\in\partial\Omega.$ (17)
By the boundary local Weyl law for the Dirichlet Laplacian paired with smooth test functions [2, 4, 8], there exist constants $A_{\Omega}>0$ and $B_{\Omega}\in\mathbb{R}$ such that
$\int_{\partial\Omega}Q_{\Lambda}(x)\,w(x)\,d\sigma(x)=A_{\Omega}\,\Lambda^{\frac{d+1}{2}}\int_{\partial\Omega}w\,d\sigma+B_{\Omega}\,\Lambda^{\frac{d}{2}}\int_{\partial\Omega}H\,w\,d\sigma+O\!\left(\Lambda^{\frac{d-1}{2}}\right),$ (18)
as $\Lambda\to\infty$, where $H$ is the mean curvature. Since $\int_{\partial\Omega}w\,d\sigma=0$, the leading term cancels. Hence there exists $C=C(\Omega,w)>0$ such that for all sufficiently large $\Lambda$,
$\left|\int_{\partial\Omega}Q_{\Lambda}(x)\,w(x)\,d\sigma(x)\right|\leq C\,\Lambda^{\frac{d}{2}}.$ (19)
Fix $k$ sufficiently large so that (19) applies to all $\Lambda\geq\lambda_{k-1}$. Choose $\Lambda$ with $\lambda_{k-1}<\Lambda<\lambda_{k}$. Then
$\int_{\partial\Omega}Q_{\Lambda}\,w\,d\sigma=\sum_{j<k}\int_{\partial\Omega}|\partial_{n}u_{j}|^{2}\,w\,d\sigma.$ (20)
Similarly, for $\Lambda^{\prime}$ with $\lambda_{k+N_{k}-1}<\Lambda^{\prime}<\lambda_{k+N_{k}}$,
$\int_{\partial\Omega}Q_{\Lambda^{\prime}}\,w\,d\sigma=\sum_{j<k+N_{k}}\int_{\partial\Omega}|\partial_{n}u_{j}|^{2}\,w\,d\sigma.$ (21)
Subtracting gives
$\sum_{j\in\mathcal{J}_{k}}\int_{\partial\Omega}|\partial_{n}u_{j}|^{2}\,w\,d\sigma=\int_{\partial\Omega}(Q_{\Lambda^{\prime}}-Q_{\Lambda})\,w\,d\sigma.$ (22)
By (19) and the triangle inequality,
$\left|\int_{\partial\Omega}(Q_{\Lambda^{\prime}}-Q_{\Lambda})\,w\,d\sigma\right|\leq\left|\int_{\partial\Omega}Q_{\Lambda^{\prime}}\,w\,d\sigma\right|+\left|\int_{\partial\Omega}Q_{\Lambda}\,w\,d\sigma\right|\leq C\Big((\Lambda^{\prime})^{\frac{d}{2}}+\Lambda^{\frac{d}{2}}\Big)\lesssim\lambda_{k+N_{k}}^{\frac{d}{2}}.$ (23)
On the other hand, by Lemma 3.1 and since $\lambda_{j}\geq\lambda_{k}$ for $j\in\mathcal{J}_{k}$,
$\sum_{j\in\mathcal{J}_{k}}E_{j}\geq c\sum_{j\in\mathcal{J}_{k}}\lambda_{j}\geq c\,N_{k}\,\lambda_{k}.$ (24)
Using Weyl asymptotics $\lambda_{k}\sim c_{\Omega}k^{2/d}$ and since $N_{k}=o(k)$, one has $\lambda_{k+N_{k}}/\lambda_{k}\to 1$. We obtain
$\frac{\left|\sum_{j\in\mathcal{J}_{k}}\int_{\partial\Omega}|\partial_{n}u_{j}|^{2}\,w\,d\sigma\right|}{\sum_{j\in\mathcal{J}_{k}}E_{j}}\lesssim\frac{\lambda_{k}^{\frac{d}{2}-1}}{N_{k}}\asymp\frac{k^{1-2/d}}{N_{k}},$ (25)
which tends to zero under the stated growth condition on $N_{k}$. This proves (15), and (16) follows by dividing by $N_{k}$. ∎
Remark 4.5 (Packet size condition).
For $d=2$ the condition $N_{k}/k^{1-2/d}\to\infty$ reduces to $N_{k}\to\infty$. For $d\geq 3$ it requires only mildly larger sublinear packets.
Remark 4.6 (Uniformity of constants).
All constants appearing in Lemma 3.1 and Proposition 4.7 are for the fixed strictly convex domain $\Omega$.
Proposition 4.7 (Packet average scale).
Let $N_{k}$ satisfy $N_{k}\to\infty$ and $N_{k}=o(k)$. Then there exist constants $c>0$ and $k_{0}$ such that for all $k\geq k_{0}$,
$\sum_{j\in\mathcal{J}_{k}}E_{j}\geq c\,N_{k}\,\lambda_{k}.$ (26)
Proof 4.8.
By Lemma 3.1 there exists $c_{0}>0$ such that $E_{j}\geq c_{0}\,\lambda_{j}$ for all $j$. Since $\lambda_{j}\geq\lambda_{k}$ for $j\in\mathcal{J}_{k}$, we obtain
$\sum_{j\in\mathcal{J}_{k}}E_{j}\geq c_{0}\,N_{k}\,\lambda_{k},$ (27)
which yields the lower bound. ∎
Remark 4.9 (Why sublinear packets).
The condition $N_{k}=o(k)$ means that the packet length is sublinear in the mode index and thus represents a short averaging window around $k$.
5 Boundary delocalization and spectral packet comparison
We now combine the packet scale lower bound with the single–mode upper bound.
Theorem 5.1 (Boundary delocalization via spectral packets).
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded $C^{\infty}$ strictly convex domain. Let $N_{k}$ satisfy $N_{k}\to\infty$ and $N_{k}=o(k)$. Then
$\frac{E_{k}}{\sum_{m=k}^{k+N_{k}-1}E_{m}}\longrightarrow 0\quad\text{as }k\to\infty.$ (28)
In particular, no single eigenmode can asymptotically carry a fixed positive proportion of the packet boundary energy on any sublinear spectral window.
Proof 5.2.
By Lemma 3.1, $E_{k}\leq C_{1}\lambda_{k}$. Proposition 4.7 gives $\sum_{m=k}^{k+N_{k}-1}E_{m}\geq c\,N_{k}\,\lambda_{k}$ for $k$ large. Hence
$\frac{E_{k}}{\sum_{m=k}^{k+N_{k}-1}E_{m}}\leq\frac{C_{1}}{c\,N_{k}}\xrightarrow[k\to\infty]{}0,$ (29)
which yields the stated limit. ∎
Corollary 5.3 (Control of weighted boundary functionals).
Let $w\in L^{\infty}(\partial\Omega)$ be any bounded geometric weight. Then the weighted energy of a single mode is dominated by the unweighted packet sum:
$\frac{\left|\int_{\partial\Omega}|\partial_{n}u_{k}|^{2}\,w(x)\,d\sigma\right|}{\sum_{m=k}^{k+N_{k}-1}E_{m}}\longrightarrow 0\quad\text{as }k\to\infty.$ (30)
Proof 5.4.
By Lemma 3.1, $|E_{k}(w)|\leq E_{k}^{abs}(w)\leq\|w\|_{\infty}E_{k}$. The result follows immediately from Theorem 5.1. ∎
Remark 5.5 (Weighted energies).
The same comparison can be extended to weighted boundary energies under additional assumptions on the weight; this is not used in the present paper. The geometric weight $w=|H-\bar{H}|$ is a natural example.
Theorem 5.6 (Absence of persistent one–sided boundary bias).
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded $C^{\infty}$ strictly convex domain. Let $w\in C^{\infty}(\partial\Omega)$ satisfy
$\int_{\partial\Omega}w\,d\sigma=0.$ (31)
Let $N_{k}\to\infty$ with $N_{k}=o(k)$ and $N_{k}/k^{1-2/d}\to\infty$. Then for every $\varepsilon>0$,
$\frac{1}{N_{k}}\#\Big\{j\in\mathcal{J}_{k}:C_{j}(w)\geq\varepsilon\Big\}\longrightarrow 0\quad\text{as }k\to\infty.$ (32)
An analogous statement holds with $C_{j}(w)\leq-\varepsilon$.
Proof 5.7.
Suppose by contradiction that there exist $\varepsilon>0$, $\delta>0$, and a subsequence such that at least $\delta N_{k}$ indices in $\mathcal{J}_{k}$ satisfy $C_{j}(w)\geq\varepsilon$. Then
$\int_{\partial\Omega}\mathcal{Q}_{k,N_{k}}(x)\,w(x)\,d\sigma(x)=\frac{1}{N_{k}}\sum_{j\in\mathcal{J}_{k}}E_{j}\,C_{j}(w)\geq\varepsilon\delta\,\frac{1}{N_{k}}\sum_{j\in\mathcal{J}_{k}}E_{j}.$ (33)
By Proposition 4.7, the right–hand side is $\gtrsim\lambda_{k}$. By Lemma 4.3,
$\int_{\partial\Omega}\mathcal{Q}_{k,N_{k}}(x)\,w(x)\,d\sigma(x)=o\!\left(\frac{1}{N_{k}}\sum_{j\in\mathcal{J}_{k}}E_{j}\right).$ (34)
By Proposition 4.7 one has $\frac{1}{N_{k}}\sum_{j\in\mathcal{J}_{k}}E_{j}\gtrsim\lambda_{k}$, hence the left–hand side is $o(\lambda_{k})$. The case $C_{j}(w)\leq-\varepsilon$ is treated analogously. This contradicts the lower bound above and finishes the proof. ∎
Corollary 5.8 (Vanishing density of strongly correlated modes).
For any fixed $w\in C^{\infty}(\partial\Omega)$ with zero mean, the set of indices $k$ such that
$\left|\int_{\partial\Omega}\rho_{k}(x)\,w(x)\,d\sigma(x)\right|\gtrsim E_{k}$ (35)
has asymptotic density zero.
Summary of the packet delocalization result
We have shown that on smooth strictly convex domains, single Dirichlet eigenfunctions cannot asymptotically dominate boundary energy relative to sublinear spectral packets. The main mode–to–packet estimate is purely static and relies only on the Rellich identity.
Conclusion
We have proved a quantitative boundary delocalization principle for Dirichlet eigenfunctions on smooth strictly convex domains, showing that no individual high–frequency eigenmode can asymptotically dominate the boundary energy of a sublinear spectral packet. The result excludes persistent boundary concentration at the level of single eigenfunctions when compared to neighboring modes in vanishing relative frequency windows.
The main mode–to–packet estimate is entirely static and relies only on the Rellich identity. For the multi–mode bias exclusion we additionally use the boundary local Weyl law; integrated boundary Weyl asymptotics for the Dirichlet Laplacian [2, 4, 8] are recorded for context. In particular, no microlocal defect measures, dynamical assumptions on the billiard flow, or ergodicity hypotheses are required. The argument is robust under eigenvalue crossings and depends only on global spectral information.
From a conceptual viewpoint, the result provides a mode–to–packet comparison mechanism on the boundary, analogous to bulk delocalization phenomena but valid in a purely deterministic and geometric setting. The sublinear packet scale reflects the minimal averaging required to control single–mode contributions using Rellich-type bounds. In addition, we obtain a multi–mode delocalization principle: no positive fraction of modes in a sublinear packet can maintain strong boundary correlation with a fixed zero–mean weight.
Several directions for further investigation suggest themselves. First, it would be natural to extend the argument to other boundary conditions or to more general elliptic operators with boundary. Second, one may ask whether analogous packet–level delocalization phenomena hold on nonconvex domains or on Riemannian manifolds with boundary, where boundary Weyl remainders are less rigid. Finally, the present comparison principle may serve as a useful input in the study of boundary spectral statistics and correlations, complementing existing microlocal and dynamical approaches.
We expect that the mode–to–packet perspective introduced here can be applied in a broader range of boundary spectral problems where individual eigenfunction control is too weak, but averaged spectral information remains accessible.

### v4

@@
 The main mode–to–packet estimate relies only on the Rellich identity.
-For the multi–mode bias exclusion we additionally use the boundary local Weyl law to obtain a packet zero–mean cancellation estimate.
+For the multi–mode bias exclusion we additionally use the boundary local Weyl law to obtain a packet zero–mean cancellation estimate at the scale $N_{k}/k^{1-2/d}\to\infty$.
 This mode–to–packet comparison is independent of eigenvalue monotonicity and is stable under eigenvalue crossings.
@@
 For general background on boundary eigenfunctions and related phenomena [14, 15].
+Classical boundary quantum ergodicity and boundary-trace completeness results include [16, 17, 18, 19].
+For recent broader developments in related semiclassical restriction and eigenfunction analysis, [20, 21].
 In contrast, the main mode–to–packet estimate is entirely static and relies only on the Rellich identity.
 For the multi–mode bias exclusion we additionally use the boundary local Weyl law to obtain packet zero–mean cancellation.
+Our approach is complementary in spirit to microlocal and dynamical studies of boundary eigenfunctions [5, 6, 12, 7, 1, 14, 15].
+We establish a deterministic mode–to–packet comparison principle based on global spectral identities: Rellich bounds for single modes and boundary local Weyl cancellation in zero–mean pairings for the multi–mode statement [2, 4, 8].
+The primary motivation for this packet–level viewpoint is structural stability under spectral deformations.
+In settings with shape perturbations or geometric flows, eigenvalue crossings and changing multiplicities are typical.
+Since the estimates below are formulated via spectral sums, they are invariant under changes of orthonormal bases inside eigenspaces and remain stable across such crossings.
 This paper shows that, on strictly convex domains, persistent boundary concentration at the level of a single mode is excluded when one compares it to a spectral packet of sublinear length.
@@
 In addition, we rule out collective boundary concentration: no positive fraction of modes inside a sublinear spectral packet can exhibit a persistent one–sided boundary bias with a fixed geometric weight of zero mean, in the spirit of averaged spectral quantities; [11, 10].
-For the multi–mode statement we assume a slightly stronger growth condition on $N_{k}$ when $d\geq 3$.
+For the multi–mode statement we assume
+$\frac{N_{k}}{k^{1-2/d}}\to\infty,$ (1)
+which is the methodological threshold for packet cancellation in the present Rellich + boundary Weyl framework.
 Theorem 1.1 (Main result - Introductory Form).
@@
 In particular, without averaging over a growing spectral window, a single eigenmode may contribute a non–negligible fraction of the packet boundary energy.
-Thus the requirement $N_{k}\to\infty$ represents the minimal averaging scale necessary to enforce mode–to–packet delocalization.
+Thus the requirement $N_{k}\to\infty$ represents the minimal averaging scale necessary to enforce single–mode to packet delocalization.
 2 Preliminaries and notation
@@
 All arguments are formulated at the level of boundary energy densities $\rho_{k}=|\partial_{n}u_{k}|^{2}$ and spectral sums.
-No ordering or differentiability of individual eigenvalues is required, and possible eigenvalue crossings do not affect the packet–level estimates.
+No ordering or differentiability of individual eigenvalues is required, possible eigenvalue crossings do not affect the packet–level estimates, and the bounds are invariant under orthonormal basis changes in eigenspaces.
 3 Rellich identity and single–mode bounds
@@
 which tends to zero under the stated growth condition on $N_{k}$. This proves (16), and (17) follows by dividing by $N_{k}$. ∎
-Remark 4.5 (Packet size condition).
-For $d=2$ the condition $N_{k}/k^{1-2/d}\to\infty$ reduces to $N_{k}\to\infty$. For $d\geq 3$ it requires only mildly larger sublinear packets.
+Proposition 4.5 (Methodological threshold for packet cancellation).
+Under the assumptions of Lemma 4.3, one has
+$\frac{\left|\sum_{j\in\mathcal{J}_{k}}\int_{\partial\Omega}|\partial_{n}u_{j}|^{2}\,w\,d\sigma\right|}{\sum_{j\in\mathcal{J}_{k}}E_{j}}\lesssim\frac{k^{1-2/d}}{N_{k}}.$ (27)
+In particular, packet-level cancellation follows whenever
+$\frac{N_{k}}{k^{1-2/d}}\to\infty.$ (28)
+If $N_{k}\lesssim k^{1-2/d}$ along a subsequence, the present method (Rellich identity combined with the boundary Weyl law) does not suffice to ensure cancellation.
+Proof 4.6.
+The estimate
+$\frac{\left|\sum_{j\in\mathcal{J}_{k}}\int_{\partial\Omega}|\partial_{n}u_{j}|^{2}\,w\,d\sigma\right|}{\sum_{j\in\mathcal{J}_{k}}E_{j}}\lesssim\frac{k^{1-2/d}}{N_{k}}$ (29)
+is exactly (26) from the proof of Lemma 4.3.
+Hence cancellation follows when $N_{k}/k^{1-2/d}\to\infty$.
+If $N_{k}\lesssim k^{1-2/d}$ along a subsequence, this upper bound is not forced to vanish, so cancellation cannot be concluded by this method alone.
+∎
+Remark 4.7 (Threshold interpretation).
+The exponent 12/d1-2/d comes from boundary Weyl fluctuations after cancellation of the principal term in the zero-mean pairing $\int_{\partial\Omega}Q_{\Lambda}w\,d\sigma$.
+For $d=2$, the condition $N_{k}/k^{1-2/d}\to\infty$ reduces to $N_{k}\to\infty$; for $d\geq 3$, it requires mildly larger sublinear packets.
 Remark 4.8 (Uniformity of constants).
-All constants appearing in Lemma 3.1 and Proposition 4.7 are for the fixed strictly convex domain $\Omega$.
+All constants appearing in Lemma 3.1 and Propositions 4.5 and 4.9 are for the fixed strictly convex domain $\Omega$.
 Proposition 4.9 (Packet average scale).
@@
 $\sum_{j\in\mathcal{J}_{k}}E_{j}\geq c\,N_{k}\,\lambda_{k}.$ (30)
-Proof 4.8.
+Proof 4.10.
 By Lemma 3.1 there exists $c_{0}>0$ such that $E_{j}\geq c_{0}\,\lambda_{j}$ for all $j$. Since $\lambda_{j}\geq\lambda_{k}$ for $j\in\mathcal{J}_{k}$, we obtain
@@
 For the multi–mode bias exclusion we additionally use the boundary local Weyl law; integrated boundary Weyl asymptotics for the Dirichlet Laplacian [2, 4, 8] are recorded for context.
-In particular, no microlocal defect measures, dynamical assumptions on the billiard flow, or ergodicity hypotheses are required.
 The argument is robust under eigenvalue crossings and depends only on global spectral information.
+This structural invariance makes the packet framework natural for deformation-type questions, where multiplicities and mode labeling may vary.
 From a conceptual viewpoint, the result provides a mode–to–packet comparison mechanism on the boundary, analogous to bulk delocalization phenomena but valid in a purely deterministic and geometric setting.
-The sublinear packet scale reflects the minimal averaging required to control single–mode contributions using Rellich-type bounds.
+For single–mode comparison, the sublinear packet regime with $N_{k}\to\infty$ and $N_{k}=o(k)$ is sufficient.
+Notably, the single–mode packet delocalization estimate is obtained without using local Weyl asymptotics.
+For packet zero–mean cancellation and multi–mode bias exclusion, the present method identifies the threshold scale $N_{k}\gg k^{1-2/d}$.
+The packet–level perspective developed here is particularly suited to averaged spectral functionals, where single–eigenfunction control is often too weak.
 In addition, we obtain a multi–mode delocalization principle: no positive fraction of modes in a sublinear packet can maintain strong boundary correlation with a fixed zero–mean weight.

### v5

@@
 Abstract
-We establish a boundary delocalization principle for high–frequency Dirichlet eigenfunctions on smooth strictly convex domains.
-The main result is an exclusion of persistent boundary concentration at the level of individual eigenmodes relative to short spectral packets of sublinear length.
-Quantitatively, we compare boundary energies of single eigenfunctions to packet sums over frequency windows of size $N_{k}=o(k)$, without asserting any asymptotic gain in magnitude.
-The main mode–to–packet estimate relies only on the Rellich identity.
-For the multi–mode bias exclusion we additionally use the boundary local Weyl law to obtain a packet zero–mean cancellation estimate at the scale $N_{k}/k^{1-2/d}\to\infty$.
-This mode–to–packet comparison is independent of eigenvalue monotonicity and is stable under eigenvalue crossings.
+We identify geometry–dependent minimal packet scales required for cancellation of boundary correlations of high–frequency Dirichlet eigenfunctions on smooth strictly convex domains.
+The main result is a threshold hierarchy: for zero–mean boundary weights, the energy–weighted packet average of boundary correlation coefficients vanishes once the packet length exceeds a scale determined by the vanishing order of curvature moments of the weight.
+In particular, the threshold $N_{k}/k^{1-2/d}\to\infty$ suffices when $\int_{\partial\Omega}w\,d\sigma=0$, while a strictly weaker threshold applies when additionally $\int_{\partial\Omega}H\,w\,d\sigma=0$, reducing in dimension $d=3$ to the minimal condition $N_{k}\to\infty$.
+The thresholds follow from the boundary local Weyl law.
+As a structural consequence of the Rellich identity alone, the single–mode share of boundary energy within any sublinear spectral packet is of order $1/N_{k}$.
+All estimates are independent of eigenvalue monotonicity and remain stable under eigenvalue crossings.
 1 Introduction
@@
 Related results on boundary delocalization and restriction estimates typically rely on microlocal analysis, defect measures, or dynamical properties of the billiard flow [5, 6, 12, 7, 1].
-For general background on boundary eigenfunctions and related phenomena [14, 15].
+For general background on boundary eigenfunctions and related phenomena, see [14, 15].
 Classical boundary quantum ergodicity and boundary-trace completeness results include [16, 17, 18, 19].
-For recent broader developments in related semiclassical restriction and eigenfunction analysis, [20, 21].
-In contrast, the main mode–to–packet estimate is entirely static and relies only on the Rellich identity.
-For the multi–mode bias exclusion we additionally use the boundary local Weyl law to obtain packet zero–mean cancellation.
+For recent broader developments in related semiclassical restriction and eigenfunction analysis, see [20, 21].
+In contrast, our approach is entirely static and relies on global spectral identities and boundary local Weyl asymptotics: the threshold hierarchy is established via the boundary local Weyl law [2, 4, 8], while the mode–to–packet comparison follows from the Rellich identity alone, with no microlocal input.
 Our approach is complementary in spirit to microlocal and dynamical studies of boundary eigenfunctions [5, 6, 12, 7, 1, 14, 15].
-We establish a deterministic mode–to–packet comparison principle based on global spectral identities: Rellich bounds for single modes and boundary local Weyl cancellation in zero–mean pairings for the multi–mode statement [2, 4, 8].
 The primary motivation for this packet–level viewpoint is structural stability under spectral deformations.
@@
 Since the estimates below are formulated via spectral sums, they are invariant under changes of orthonormal bases inside eigenspaces and remain stable across such crossings.
-This paper shows that, on strictly convex domains, persistent boundary concentration at the level of a single mode is excluded when one compares it to a spectral packet of sublinear length.
-We establish a packet–level delocalization principle: for any sublinear window $N_{k}=o(k)$, the ratio between a single boundary energy and the packet sum tends to zero.
-The proof uses the Rellich identity.
-In addition, we rule out collective boundary concentration: no positive fraction of modes inside a sublinear spectral packet can exhibit a persistent one–sided boundary bias with a fixed geometric weight of zero mean, in the spirit of averaged spectral quantities; [11, 10].
-For the multi–mode statement we assume
-$\frac{N_{k}}{k^{1-2/d}}\to\infty,$ (1)
-which is the methodological threshold for packet cancellation in the present Rellich + boundary Weyl framework.
-Theorem 1.1 (Main result - Introductory Form).
+Moreover, all constants are uniform over compact $C^{\infty}$ families of strictly convex domains (see Remark 5.9), making the threshold hierarchy directly applicable in dynamical settings where the domain evolves continuously and cancellation must be controlled uniformly in time.
+The principal new contribution of this paper is the identification of geometry–dependent minimal packet scales for boundary correlation cancellation on strictly convex domains.
+For any zero–mean weight $w\in C^{\infty}(\partial\Omega)$, the energy–weighted packet average of boundary correlation coefficients vanishes once $N_{k}$ crosses a threshold determined by the vanishing order of curvature moments of $w$, in the spirit of averaged spectral asymptotics [11, 10].
+This threshold forms a hierarchy of two levels: the first, $N_{k}/k^{1-2/d}\to\infty$, suffices when $\int_{\partial\Omega}w\,d\sigma=0$; the second, strictly weaker level $N_{k}/k^{(d-3)/d}\to\infty$ applies when additionally $\int_{\partial\Omega}H\,w\,d\sigma=0$, reducing for $d=3$ to the minimal condition $N_{k}\to\infty$.
+The hierarchy is stratified by the vanishing order of spectral moments of $w$ with respect to boundary geometry and is a new structural feature of the boundary spectral problem with no analogue in bulk delocalization results.
+We emphasize that the novelty lies not in the boundary Weyl expansion itself, which is classical [2, 4, 8], but in the extraction of explicit minimal packet-length scales forced by successive moment cancellations and the resulting detectability hierarchy: the question ”how much spectral averaging is needed to make boundary correlations undetectable by the Rellich–Weyl mechanism?” is answered quantitatively and stratified by the geometry of $w$.
+Theorem 1.1 (Threshold hierarchy for boundary cancellation — Introductory Form).
+Let $\Omega\subset\mathbb{R}^{d}$ be a bounded $C^{\infty}$ strictly convex domain, $d\geq 2$, and let $\{(\lambda_{k},u_{k})\}_{k\geq 1}$ be the Dirichlet eigenpairs of $\Delta_{\Omega}$.
+Let $w\in C^{\infty}(\partial\Omega)$ and let $N_{k}$ satisfy $N_{k}\to\infty$ and $N_{k}=o(k)$.
+  1. (i)
+    If $\int_{\partial\Omega}w\,d\sigma=0$ and $N_{k}/k^{1-2/d}\to\infty$, then
+    $\frac{\displaystyle\sum_{j=k}^{k+N_{k}-1}\int_{\partial\Omega}|\partial_{n}u_{j}|^{2}\,w\,d\sigma}{\displaystyle\sum_{j=k}^{k+N_{k}-1}E_{j}}\longrightarrow 0\quad\text{as }k\to\infty.$ (1)
+  2. (ii)
+    If additionally $\int_{\partial\Omega}H\,w\,d\sigma=0$ and $N_{k}/k^{(d-3)/d}\to\infty$, the same conclusion holds. For $d=3$, case (ii) holds under the minimal condition $N_{k}\to\infty$.
+These thresholds are geometry–dependent: they are determined by the vanishing order of curvature moments of $w$ with respect to the boundary of $\Omega$.
+We next record a structural comparison result following directly from the Rellich identity, which explains the intrinsic $1/N_{k}$ scaling of single–mode contributions within spectral packets and does not require Weyl asymptotics or microlocal analysis.
+Theorem 1.2 (Structural consequence: mode-to-packet comparison — Introductory Form).
 Let $\Omega\subset\mathbb{R}^{d}$ be a bounded $C^{\infty}$ strictly convex domain, $d\geq 2$, and let $\{(\lambda_{k},u_{k})\}_{k\geq 1}$ be the Dirichlet eigenpairs of $\Delta_{\Omega}$.
 Let $N_{k}$ be any sublinear spectral window with $N_{k}\to\infty$ and $N_{k}=o(k)$.
-Then
-$\frac{E_{k}}{\sum_{m=k}^{k+N_{k}-1}E_{m}}\longrightarrow 0\quad\text{as }k\to\infty.$ (2)
-In particular, no individual high–frequency eigenfunction can asymptotically carry a fixed positive proportion of the boundary energy of a sublinear spectral packet.
-Remark 1.2 (Sharpness of the packet length condition).
-The condition $N_{k}\to\infty$ is essential at the level of the present method. Indeed, for a fixed packet length $N$, the Rellich identity and the Weyl law alone do not imply that
-$\frac{E_{k}}{\sum_{m=k}^{k+N-1}E_{m}}\longrightarrow 0.$ (3)
+Then there exist constants $0<c\leq C$ depending only on $\Omega$ such that
+$\frac{c}{N_{k}}\;\leq\;\frac{E_{k}}{\sum_{m=k}^{k+N_{k}-1}E_{m}}\;\leq\;\frac{C}{N_{k}}\qquad\text{for all sufficiently large }k.$ (2)
+In particular, no individual high–frequency eigenfunction can asymptotically carry a fixed positive proportion of the boundary energy of a sublinear spectral packet, and the single–mode share decays at exactly the rate $1/N_{k}$.
+Remark 1.3 (Sharpness of the rate $1/N_{k}$).
+The two–sided bound $c/N_{k}\leq E_{k}/\sum E_{m}\leq C/N_{k}$ shows that the rate $1/N_{k}$ is optimal: it cannot be replaced by a faster-decaying sequence in general.
+For a fixed packet length $N$, the ratio satisfies
+$\frac{c}{N}\;\leq\;\frac{E_{k}}{\sum_{m=k}^{k+N-1}E_{m}}\;\leq\;\frac{C}{N},$ (3)
+which is bounded away from zero.
 In particular, without averaging over a growing spectral window, a single eigenmode may contribute a non–negligible fraction of the packet boundary energy.
-Thus the requirement $N_{k}\to\infty$ represents the minimal averaging scale necessary to enforce single–mode to packet delocalization.
+The condition $N_{k}\to\infty$ is both necessary and sufficient for the ratio to tend to zero, and $1/N_{k}$ is the precise rate of decay.
 2 Preliminaries and notation
@@
 $E_{k}(w)$ $=\int_{\partial\Omega}\rho_{k}(x)\,w(x)\,d\sigma(x),$ (8)
-$E_{k}^{abs}(w)$ $=\int_{\partial\Omega}\rho_{k}(x)\,|w(x)|\,d\sigma(x),$
-$\mathcal{E}_{k,N_{k}}(w)$ $=\frac{1}{N_{k}}\sum_{j\in\mathcal{J}_{k}}E_{j}^{abs}(w).$
+$E_{k}^{abs}(w)$ $=\int_{\partial\Omega}\rho_{k}(x)\,|w(x)|\,d\sigma(x).$
 Definition 2.2 (Boundary correlation coefficient).
@@
 3 Rellich identity and single–mode bounds
-The Rellich identity yields the correct scaling for the boundary energy of a single eigenfunction on strictly convex domains; [13, 3].
+The Rellich identity yields the correct scaling for the boundary energy of a single eigenfunction on strictly convex domains; [13].
 Lemma 3.1 (Single–mode boundary energy bound).
@@
 which gives the claimed bounds. The weighted estimate follows from $|E_{k}(w)|\leq E_{k}^{abs}(w)\leq\|w\|_{L^{\infty}}E_{k}$. ∎
-4 Packet scales and a contextual Weyl remark
+Remark 3.3 (Role of strict convexity and smoothness).
+Strict convexity enters the Rellich argument in an essential way: it ensures that $g(x)=\langle x-x_{0},\nu(x)\rangle>0$ uniformly on $\partial\Omega$ for any interior point $x_{0}$, which is precisely the condition needed to obtain the two-sided bounds $c\lambda_{k}\leq E_{k}\leq C\lambda_{k}$.
+On a merely star-shaped domain one obtains only one-sided upper control.
+Smoothness of $\partial\Omega$ is needed separately for the boundary local Weyl expansion (22): the classical $O(\Lambda^{(d-1)/2})$ remainder in the integrated pairing with smooth $w$ is available on any $C^{\infty}$ domain [2, 4], without non-periodicity or dynamical assumptions on the billiard flow.
+4 Packet energy scales and cancellation thresholds
 The integrated boundary energy admits a Weyl-type leading asymptotic on smooth strictly convex domains; [2, 4] We record this only for context.
@@
 as $\Lambda\to\infty$; [2, 4, 9]. This is recorded for context and is not used in the proof of the main theorem.
+Proposition 4.3 (Packet average scale).
+Let $N_{k}$ satisfy $N_{k}\to\infty$ and $N_{k}=o(k)$. Then there exist constants $0<c\leq C$ and $k_{0}$ such that for all $k\geq k_{0}$,
+$c\,N_{k}\,\lambda_{k}\;\leq\;\sum_{j\in\mathcal{J}_{k}}E_{j}\;\leq\;C\,N_{k}\,\lambda_{k}.$ (16)
+Proof 4.4.
+Lower bound. By Lemma 3.1 there exists $c_{0}>0$ such that $E_{j}\geq c_{0}\,\lambda_{j}$ for all $j$. Since $\lambda_{j}\geq\lambda_{k}$ for $j\in\mathcal{J}_{k}$,
+$\sum_{j\in\mathcal{J}_{k}}E_{j}\geq c_{0}\,N_{k}\,\lambda_{k}.$ (17)
+Upper bound. By Lemma 3.1 there exists $C_{0}>0$ such that $E_{j}\leq C_{0}\,\lambda_{j}$ for all $j$. For $j\in\mathcal{J}_{k}$ one has $\lambda_{j}\leq\lambda_{k+N_{k}-1}$, so
+$\sum_{j\in\mathcal{J}_{k}}E_{j}\leq C_{0}\,N_{k}\,\lambda_{k+N_{k}-1}.$ (18)
+By Weyl asymptotics $\lambda_{k}\sim c_{\Omega}k^{2/d}$ and $N_{k}=o(k)$, one has $\lambda_{k+N_{k}-1}/\lambda_{k}\to 1$, so $\lambda_{k+N_{k}-1}\leq 2\lambda_{k}$ for all sufficiently large $k$.
+Hence $\sum_{j\in\mathcal{J}_{k}}E_{j}\leq 2C_{0}\,N_{k}\,\lambda_{k}$ for $k\geq k_{0}$.
+∎
+Remark 4.5 (Why sublinear packets).
+The condition $N_{k}=o(k)$ means that the packet length is sublinear in the mode index and thus represents a short averaging window around $k$.
 Lemma 4.6 (Packet zero–mean cancellation).
@@
 $\int_{\partial\Omega}\mathcal{Q}_{k,N_{k}}(x)\,w(x)\,d\sigma(x)=o\!\left(\frac{1}{N_{k}}\sum_{j\in\mathcal{J}_{k}}E_{j}\right)\qquad(k\to\infty).$ (20)
-Proof 4.4.
+Proof 4.7.
 For $\Lambda>0$ define the boundary spectral density
@@
 $\int_{\partial\Omega}Q_{\Lambda}(x)\,w(x)\,d\sigma(x)=A_{\Omega}\,\Lambda^{\frac{d+1}{2}}\int_{\partial\Omega}w\,d\sigma+B_{\Omega}\,\Lambda^{\frac{d}{2}}\int_{\partial\Omega}H\,w\,d\sigma+O\!\left(\Lambda^{\frac{d-1}{2}}\right),$ (22)
-as $\Lambda\to\infty$, where $H$ is the mean curvature.
+as $\Lambda\to\infty$.
+The two-term structure of (22) follows from the boundary heat kernel expansion [8]: the first term arises because the leading heat kernel coefficient on the boundary is constant (the flat-space contribution), so its pairing with $w$ gives $\Lambda^{(d+1)/2}\!\int w\,d\sigma$; the second term arises because the next coefficient is proportional to the mean curvature $H$ [8], so its pairing with $w$ gives $\Lambda^{d/2}\!\int Hw\,d\sigma$.
+The remainder $O(\Lambda^{(d-1)/2})$ corresponds to the $O(\Lambda^{(d-2)/2})$ error in the standard Weyl expansion after integration.
 Since $\int_{\partial\Omega}w\,d\sigma=0$, the leading term cancels.
@@
 which tends to zero under the stated growth condition on $N_{k}$. This proves (19), and (20) follows by dividing by $N_{k}$. ∎
-Proposition 4.5 (Methodological threshold for packet cancellation).
+Extracting the quantitative rate from the above proof yields the following detectability threshold, which makes explicit the minimal packet scale forced by the Weyl remainder.
+Proposition 4.8 (Method-detectability threshold from Weyl remainder bounds).
 Under the assumptions of Lemma 4.6, one has
 $\frac{\left|\sum_{j\in\mathcal{J}_{k}}\int_{\partial\Omega}|\partial_{n}u_{j}|^{2}\,w\,d\sigma\right|}{\sum_{j\in\mathcal{J}_{k}}E_{j}}\lesssim\frac{k^{1-2/d}}{N_{k}}.$ (30)
-In particular, packet-level cancellation follows whenever
+In particular, packet-level cancellation is guaranteed whenever
 $\frac{N_{k}}{k^{1-2/d}}\to\infty.$ (31)
-If $N_{k}\lesssim k^{1-2/d}$ along a subsequence, the present method (Rellich identity combined with the boundary Weyl law) does not suffice to ensure cancellation.
-Proof 4.6.
+If $N_{k}\lesssim k^{1-2/d}$ along a subsequence, the present method (Rellich identity combined with the boundary Weyl law) does not suffice to detect cancellation.
+Proof 4.9.
 The estimate
@@
 ∎
-Remark 4.7 (Threshold interpretation).
-The exponent 12/d1-2/d comes from boundary Weyl fluctuations after cancellation of the principal term in the zero-mean pairing $\int_{\partial\Omega}Q_{\Lambda}w\,d\sigma$.
-For $d=2$, the condition $N_{k}/k^{1-2/d}\to\infty$ reduces to $N_{k}\to\infty$; for $d\geq 3$, it requires mildly larger sublinear packets.
+Remark 4.10 (Method-detectability and openness of sharpness).
+The threshold $N_{k}\gg k^{1-2/d}$ is a sufficient condition forced by the $O(\Lambda^{d/2})$ remainder in the boundary local Weyl expansion after cancellation of the leading term via $\int_{\partial\Omega}w\,d\sigma=0$.
+The exponent 12/d1-2/d reflects the size of second-term Weyl fluctuations: for $d=2$ the condition reduces to $N_{k}\to\infty$; for $d\geq 3$ it requires mildly larger sublinear packets.
+The question of whether cancellation can occur below this scale—i.e., whether $N_{k}\gg k^{1-2/d}$ is also necessary—depends on finer spectral information beyond the classical Weyl remainder bounds and remains open.
+The present threshold is, however, the minimal scale at which the Rellich–Weyl mechanism can detect cancellation.
+It is therefore the natural packet size for deformation-stable arguments in which cancellation must be uniform over compact families of domains.
+This makes the packet-scale threshold particularly natural in geometric deformation settings (e.g.
+smooth domain flows), where modewise cancellation must persist uniformly in time.
+We use only the classical $O(\Lambda^{(d-1)/2})$ remainder in the integrated pairing with smooth $w$, available on any smooth strictly convex domain without dynamical or non-periodicity assumptions; sharper remainders are not required.
 Remark 4.11 (Uniformity of constants).
-All constants appearing in Lemma 3.1 and Propositions 4.5 and 4.9 are for the fixed strictly convex domain $\Omega$.
-Proposition 4.9 (Packet average scale).
-Let $N_{k}$ satisfy $N_{k}\to\infty$ and $N_{k}=o(k)$. Then there exist constants $c>0$ and $k_{0}$ such that for all $k\geq k_{0}$,
-$\sum_{j\in\mathcal{J}_{k}}E_{j}\geq c\,N_{k}\,\lambda_{k}.$ (30)
-Proof 4.10.
-By Lemma 3.1 there exists $c_{0}>0$ such that $E_{j}\geq c_{0}\,\lambda_{j}$ for all $j$. Since $\lambda_{j}\geq\lambda_{k}$ for $j\in\mathcal{J}_{k}$, we obtain
-$\sum_{j\in\mathcal{J}_{k}}E_{j}\geq c_{0}\,N_{k}\,\lambda_{k},$ (31)
-which yields the lower bound. ∎
-Remark 4.11 (Why sublinear packets).
-The condition $N_{k}=o(k)$ means that the packet length is sublinear in the mode index and thus represents a short averaging window around $k$.
-5 Boundary delocalization and spectral packet comparison
-We now combine the packet scale lower bound with the single–mode upper bound.
-Theorem 5.1 (Boundary delocalization via spectral packets).
-Let $\Omega\subset\mathbb{R}^{d}$ be a bounded $C^{\infty}$ strictly convex domain. Let $N_{k}$ satisfy $N_{k}\to\infty$ and $N_{k}=o(k)$. Then
-$\frac{E_{k}}{\sum_{m=k}^{k+N_{k}-1}E_{m}}\longrightarrow 0\quad\text{as }k\to\infty.$ (32)
-In particular, no single eigenmode can asymptotically carry a fixed positive proportion of the packet boundary energy on any sublinear spectral window.
+All constants appearing in Lemma 3.1 and Propositions 4.8 and 4.3 are for the fixed strictly convex domain $\Omega$.
+Proposition 4.12 (Enhanced cancellation under double vanishing moment).
+Let $\Omega\subset\mathbb{R}^{d}$ be a bounded $C^{\infty}$ strictly convex domain, $d\geq 2$, and let $H$ denote the mean curvature of $\partial\Omega$. Let $w\in C^{\infty}(\partial\Omega)$ satisfy
+$\int_{\partial\Omega}w\,d\sigma=0\quad\text{and}\quad\int_{\partial\Omega}H\,w\,d\sigma=0.$ (33)
+Let $N_{k}\to\infty$ with $N_{k}=o(k)$ and $N_{k}/k^{(d-3)/d}\to\infty$, and set $\mathcal{J}_{k}=\{k,\dots,k+N_{k}-1\}$. Then
+$\sum_{j\in\mathcal{J}_{k}}\int_{\partial\Omega}|\partial_{n}u_{j}|^{2}\,w\,d\sigma=o\!\left(\sum_{j\in\mathcal{J}_{k}}E_{j}\right)\qquad(k\to\infty).$ (34)
+This threshold is strictly weaker than $N_{k}/k^{1-2/d}\to\infty$ for all $d\geq 3$.
+Proof 4.13.
+Since both $\int_{\partial\Omega}w\,d\sigma=0$ and $\int_{\partial\Omega}H\,w\,d\sigma=0$, both leading terms in the boundary local Weyl expansion (22) cancel, yielding
+$\left|\int_{\partial\Omega}Q_{\Lambda}(x)\,w(x)\,d\sigma(x)\right|\leq C\,\Lambda^{\frac{d-1}{2}}$ (35)
+for all sufficiently large $\Lambda$. Applying the same telescoping argument as in the proof of Lemma 4.6 and using the two-sided bound of Proposition 4.3, one obtains
+$\frac{\left|\sum_{j\in\mathcal{J}_{k}}\int_{\partial\Omega}|\partial_{n}u_{j}|^{2}\,w\,d\sigma\right|}{\sum_{j\in\mathcal{J}_{k}}E_{j}}\;\lesssim\;\frac{\lambda_{k}^{\frac{d-3}{2}}}{N_{k}}\;\asymp\;\frac{k^{\frac{d-3}{d}}}{N_{k}},$ (36)
+which tends to zero under $N_{k}/k^{(d-3)/d}\to\infty$.
+The comparison with the first-moment threshold follows from (d3)/d<12/d(d-3)/d<1-2/d for all $d\geq 2$, with strict improvement for $d\geq 3$: (i) $d=2$: both thresholds reduce to $N_{k}\to\infty$; (ii) $d=3$: first-moment threshold is $N_{k}\gg k^{1/3}$, double-moment threshold reduces to $N_{k}\to\infty$; (iii) $d\geq 4$: first-moment threshold is $N_{k}\gg k^{1-2/d}$, double-moment threshold is $N_{k}\gg k^{(d-3)/d}$ with (d3)/d<12/d(d-3)/d<1-2/d.
+∎
+5 Cancellation hierarchy and mode-to-packet comparison
+We now establish the threshold hierarchy for boundary cancellation and derive the structural mode-to-packet comparison as a consequence of the Rellich identity.
+Theorem 5.1 (Sharp two-sided packet energy comparison).
+Let $\Omega\subset\mathbb{R}^{d}$ be a bounded $C^{\infty}$ strictly convex domain.
+Let $N_{k}$ satisfy $N_{k}\to\infty$ and $N_{k}=o(k)$.
+Then there exist constants $0<c\leq C$ depending only on $\Omega$ and $k_{0}$ such that for all $k\geq k_{0}$,
+$\frac{c}{N_{k}}\;\leq\;\frac{E_{k}}{\displaystyle\sum_{m=k}^{k+N_{k}-1}E_{m}}\;\leq\;\frac{C}{N_{k}}.$ (37)
+In particular, no single eigenmode can asymptotically carry a fixed positive proportion of the packet boundary energy on any sublinear spectral window, and the rate of decay $1/N_{k}$ is sharp.
 Proof 5.2.
-By Lemma 3.1, $E_{k}\leq C_{1}\lambda_{k}$. Proposition 4.9 gives $\sum_{m=k}^{k+N_{k}-1}E_{m}\geq c\,N_{k}\,\lambda_{k}$ for $k$ large. Hence
-$\frac{E_{k}}{\sum_{m=k}^{k+N_{k}-1}E_{m}}\leq\frac{C_{1}}{c\,N_{k}}\xrightarrow[k\to\infty]{}0,$ (33)
-which yields the stated limit. ∎
+Upper bound. By Lemma 3.1, $E_{k}\leq C_{1}\lambda_{k}$. By Proposition 4.3, $\sum_{m=k}^{k+N_{k}-1}E_{m}\geq c_{1}\,N_{k}\,\lambda_{k}$ for $k\geq k_{0}$. Hence
+$\frac{E_{k}}{\sum_{m=k}^{k+N_{k}-1}E_{m}}\leq\frac{C_{1}}{c_{1}\,N_{k}}.$ (38)
+Lower bound. By Lemma 3.1, $E_{k}\geq c_{0}\lambda_{k}$. By the upper bound in Proposition 4.3, $\sum_{m=k}^{k+N_{k}-1}E_{m}\leq C_{2}\,N_{k}\,\lambda_{k}$ for $k\geq k_{0}$. Hence
+$\frac{E_{k}}{\sum_{m=k}^{k+N_{k}-1}E_{m}}\geq\frac{c_{0}}{C_{2}\,N_{k}}.$ (39)
+Setting $c=c_{0}/C_{2}$ and $C=C_{1}/c_{1}$ completes the proof. ∎
+Remark 5.3 (Optimality of the rate).
+The two-sided bound shows that $E_{k}/\sum E_{m}\asymp 1/N_{k}$ precisely.
+The lower bound is not a consequence of any anomalous concentration: it holds for every domain and every sequence $N_{k}=o(k)$, simply because $E_{k}\geq c\lambda_{k}$ and $\sum E_{m}\leq CN_{k}\lambda_{k}$ both follow from the Rellich identity alone.
+In other words, the packet sum cannot grow faster than $N_{k}\lambda_{k}$, so the single-mode contribution can never be smaller than order $1/N_{k}$ relative to the packet.
+This demonstrates that the delocalization rate is dictated by the global spectral structure, not by any special geometric property of individual modes.
 Corollary 5.4 (Control of weighted boundary functionals).
@@
 $\frac{\left|\int_{\partial\Omega}|\partial_{n}u_{k}|^{2}\,w(x)\,d\sigma\right|}{\sum_{m=k}^{k+N_{k}-1}E_{m}}\longrightarrow 0\quad\text{as }k\to\infty.$ (40)
-Proof 5.4.
+Proof $5.$
 By Lemma 3.1, $|E_{k}(w)|\leq E_{k}^{abs}(w)\leq\|w\|_{\infty}E_{k}$. The result follows immediately from Theorem 5.1. ∎
@@
 The geometric weight $w=|H-\bar{H}|$ is a natural example.
-Theorem 5.6 (Absence of persistent one–sided boundary bias).
+Theorem 5.7 (Packet zero–mean cancellation: multi–mode form).
 Let $\Omega\subset\mathbb{R}^{d}$ be a bounded $C^{\infty}$ strictly convex domain. Let $w\in C^{\infty}(\partial\Omega)$ satisfy
 $\int_{\partial\Omega}w\,d\sigma=0.$ (41)
-Let $N_{k}\to\infty$ with $N_{k}=o(k)$ and $N_{k}/k^{1-2/d}\to\infty$. Then for every $\varepsilon>0$,
-$\frac{1}{N_{k}}\#\Big\{j\in\mathcal{J}_{k}:C_{j}(w)\geq\varepsilon\Big\}\longrightarrow 0\quad\text{as }k\to\infty.$ (36)
-An analogous statement holds with $C_{j}(w)\leq-\varepsilon$.
-Proof 5.7.
-Suppose by contradiction that there exist $\varepsilon>0$, $\delta>0$, and a subsequence such that at least $\delta N_{k}$ indices in $\mathcal{J}_{k}$ satisfy $C_{j}(w)\geq\varepsilon$. Then
-$\int_{\partial\Omega}\mathcal{Q}_{k,N_{k}}(x)\,w(x)\,d\sigma(x)=\frac{1}{N_{k}}\sum_{j\in\mathcal{J}_{k}}E_{j}\,C_{j}(w)\geq\varepsilon\delta\,\frac{1}{N_{k}}\sum_{j\in\mathcal{J}_{k}}E_{j}.$ (37)
-By Proposition 4.9, the right–hand side is $\gtrsim\lambda_{k}$. By Lemma 4.3,
-$\int_{\partial\Omega}\mathcal{Q}_{k,N_{k}}(x)\,w(x)\,d\sigma(x)=o\!\left(\frac{1}{N_{k}}\sum_{j\in\mathcal{J}_{k}}E_{j}\right).$ (38)
-By Proposition 4.9 one has $\frac{1}{N_{k}}\sum_{j\in\mathcal{J}_{k}}E_{j}\gtrsim\lambda_{k}$, hence the left–hand side is $o(\lambda_{k})$.
-The case $C_{j}(w)\leq-\varepsilon$ is treated analogously.
-This contradicts the lower bound above and finishes the proof.
-∎
-Corollary 5.8 (Vanishing density of strongly correlated modes).
-For any fixed $w\in C^{\infty}(\partial\Omega)$ with zero mean, the set of indices $k$ such that
-$\left|\int_{\partial\Omega}\rho_{k}(x)\,w(x)\,d\sigma(x)\right|\gtrsim E_{k}$ (39)
-has asymptotic density zero.
-Summary of the packet delocalization result
-We have shown that on smooth strictly convex domains, single Dirichlet eigenfunctions cannot asymptotically dominate boundary energy relative to sublinear spectral packets.
-The main mode–to–packet estimate is purely static and relies only on the Rellich identity.
+Let $N_{k}\to\infty$ with $N_{k}=o(k)$ and $N_{k}/k^{1-2/d}\to\infty$. Then the energy–weighted average of the boundary correlation coefficients within the packet vanishes:
+$\frac{\displaystyle\sum_{j\in\mathcal{J}_{k}}E_{j}\,C_{j}(w)}{\displaystyle\sum_{j\in\mathcal{J}_{k}}E_{j}}\longrightarrow 0\quad\text{as }k\to\infty.$ (42)
+Equivalently,
+$\frac{\displaystyle\sum_{j\in\mathcal{J}_{k}}\int_{\partial\Omega}|\partial_{n}u_{j}|^{2}\,w\,d\sigma}{\displaystyle\sum_{j\in\mathcal{J}_{k}}E_{j}}\longrightarrow 0.$ (43)
+Proof 5.8.
+By Lemma 4.6,
+$\sum_{j\in\mathcal{J}_{k}}\int_{\partial\Omega}|\partial_{n}u_{j}|^{2}\,w\,d\sigma=o\!\left(\sum_{j\in\mathcal{J}_{k}}E_{j}\right).$ (44)
+Dividing by $\sum_{j\in\mathcal{J}_{k}}E_{j}$ (which is $\asymp N_{k}\lambda_{k}$ by Proposition 4.3) gives the result. ∎
+Remark 5.9 (Quantitative rate and uniformity over compact domain families).
+The proof of Theorem 5.7 yields the explicit quantitative bound
+$\left|\frac{\displaystyle\sum_{j\in\mathcal{J}_{k}}\int_{\partial\Omega}|\partial_{n}u_{j}|^{2}\,w\,d\sigma}{\displaystyle\sum_{j\in\mathcal{J}_{k}}E_{j}}\right|\leq C(\Omega,w)\,\frac{k^{1-2/d}}{N_{k}},$ (45)
+where $C(\Omega,w)$ depends on $\Omega$ only through its $C^{\infty}$ geometry (Rellich constants and the boundary Weyl law remainder) and on $w$ only through $\|w\|_{C^{\infty}(\partial\Omega)}$.
+Since the boundary Weyl law remainder bound (23) is uniform on compact $C^{\infty}$ families of strictly convex domains [2, 4], and the Rellich lower bound for $\sum_{j\in\mathcal{J}_{k}}E_{j}$ is similarly uniform by compactness, the constant $C(\Omega,w)$ may be chosen uniformly whenever $\Omega$ varies in a compact $C^{\infty}$ family and $\|w\|_{C^{\infty}}$ is uniformly bounded.
+In particular, for a compact smooth flow $\{\Omega_{t}\}_{t\in[0,T]}$ with weight $w_{t}$ satisfying $\int_{\partial\Omega_{t}}w_{t}\,d\sigma_{t}=0$ and $\sup_{t}\|w_{t}\|_{C^{\infty}}<\infty$,
+$\sup_{t\in[0,T]}\left|\frac{\displaystyle\sum_{j\in\mathcal{J}_{k}}\int_{\partial\Omega_{t}}|\partial_{n}u_{j}(t)|^{2}\,w_{t}\,d\sigma_{t}}{\displaystyle\sum_{j\in\mathcal{J}_{k}}E_{j}(t)}\right|\leq C\,\frac{k^{1-2/d}}{N_{k}}\longrightarrow 0,$ (46)
+with $C$ independent of $t$.
+Theorem 5.10 (Enhanced packet cancellation under double vanishing moment).
+Let $\Omega\subset\mathbb{R}^{d}$ be a bounded $C^{\infty}$ strictly convex domain. Let $w\in C^{\infty}(\partial\Omega)$ satisfy
+$\int_{\partial\Omega}w\,d\sigma=0\quad\text{and}\quad\int_{\partial\Omega}H\,w\,d\sigma=0.$ (47)
+Let $N_{k}\to\infty$ with $N_{k}=o(k)$ and $N_{k}/k^{(d-3)/d}\to\infty$. Then
+$\frac{\displaystyle\sum_{j\in\mathcal{J}_{k}}E_{j}\,C_{j}(w)}{\displaystyle\sum_{j\in\mathcal{J}_{k}}E_{j}}\longrightarrow 0\quad\text{as }k\to\infty.$ (48)
+For $d=3$ this holds under the minimal condition $N_{k}\to\infty$, and for $d\geq 4$ the threshold $N_{k}\gg k^{(d-3)/d}$ is strictly weaker than the threshold $N_{k}\gg k^{1-2/d}$ of Theorem 5.7.
+Proof 5.11.
+By Proposition 4.12,
+$\sum_{j\in\mathcal{J}_{k}}\int_{\partial\Omega}|\partial_{n}u_{j}|^{2}\,w\,d\sigma=o\!\left(\sum_{j\in\mathcal{J}_{k}}E_{j}\right).$ (49)
+Dividing by $\sum_{j\in\mathcal{J}_{k}}E_{j}$ gives the result. ∎
+Remark 5.12 (Two-level hierarchy and extensibility).
+Theorems 5.7 and 5.10 together exhibit a two-level hierarchy of cancellation thresholds indexed by the vanishing order of boundary curvature moments of $w$: (i) $\int w\,d\sigma=0$ only: threshold $N_{k}\gg k^{1-2/d}$; (ii) $\int w\,d\sigma=0$ and $\int H\,w\,d\sigma=0$: threshold $N_{k}\gg k^{(d-3)/d}$, strictly weaker for $d\geq 3$.
+These two levels are canonical: they correspond precisely to the first two universal terms in the Safarov–Vassiliev boundary Weyl expansion (22), and each level is determined by the vanishing order of the corresponding expansion coefficient against $w$.
+In principle the same argument extends to higher-order moments whenever further terms in the Weyl expansion are available and vanish against $w$; we restrict to the first two levels since they cover the main geometric cases and are sufficient for applications.
+The natural weight $w=H-\bar{H}$ satisfies the first condition by definition, but not necessarily the second unless $H$ is constant (i.e., $\Omega$ is a ball).
+For perturbations away from a ball, both theorems apply with the appropriate threshold, providing finer cancellation-scale information near round domains.
 Conclusion
-We have proved a quantitative boundary delocalization principle for Dirichlet eigenfunctions on smooth strictly convex domains, showing that no individual high–frequency eigenmode can asymptotically dominate the boundary energy of a sublinear spectral packet.
-The result excludes persistent boundary concentration at the level of single eigenfunctions when compared to neighboring modes in vanishing relative frequency windows.
-The main mode–to–packet estimate is entirely static and relies only on the Rellich identity.
-For the multi–mode bias exclusion we additionally use the boundary local Weyl law; integrated boundary Weyl asymptotics for the Dirichlet Laplacian [2, 4, 8] are recorded for context.
+The principal new contribution of this paper is the identification of a threshold hierarchy for boundary correlation cancellation of Dirichlet eigenfunctions on smooth strictly convex domains.
+The minimal packet scale required for the energy–weighted average of boundary correlation coefficients to vanish is determined by the vanishing order of curvature moments of the weight with respect to domain geometry.
+Theorems 5.7 and 5.10 establish two levels of this hierarchy, via the boundary local Weyl law [2, 4, 8]: (i) If $\int_{\partial\Omega}w\,d\sigma=0$, packet zero–mean cancellation holds at threshold $N_{k}\gg k^{1-2/d}$.
+(ii) If additionally $\int_{\partial\Omega}H\,w\,d\sigma=0$, the threshold weakens to $N_{k}\gg k^{(d-3)/d}$, which for $d=3$ reduces to $N_{k}\to\infty$.
+This hierarchy, stratified by the vanishing order of boundary curvature moments of $w$, is a new structural feature of the boundary spectral problem with no analogue in bulk delocalization results.
+As a structural consequence of the Rellich identity alone, Theorem 5.1 establishes that the ratio of the boundary energy of a single eigenmode to the packet sum satisfies $c/N_{k}\leq E_{k}/\sum_{m=k}^{k+N_{k}-1}E_{m}\leq C/N_{k}$.
+The rate $1/N_{k}$ is sharp and follows from global Rellich bounds alone, without any use of Weyl asymptotics or microlocal analysis.
 The argument is robust under eigenvalue crossings and depends only on global spectral information.
-This structural invariance makes the packet framework natural for deformation-type questions, where multiplicities and mode labeling may vary.
-From a conceptual viewpoint, the result provides a mode–to–packet comparison mechanism on the boundary, analogous to bulk delocalization phenomena but valid in a purely deterministic and geometric setting.
-For single–mode comparison, the sublinear packet regime with $N_{k}\to\infty$ and $N_{k}=o(k)$ is sufficient.
-Notably, the single–mode packet delocalization estimate is obtained without using local Weyl asymptotics.
-For packet zero–mean cancellation and multi–mode bias exclusion, the present method identifies the threshold scale $N_{k}\gg k^{1-2/d}$.
-The packet–level perspective developed here is particularly suited to averaged spectral functionals, where single–eigenfunction control is often too weak.
-In addition, we obtain a multi–mode delocalization principle: no positive fraction of modes in a sublinear packet can maintain strong boundary correlation with a fixed zero–mean weight.
+This structural invariance makes the packet framework particularly natural for deformation–type questions, where multiplicities and mode labeling may vary.
 Several directions for further investigation suggest themselves.
@@
 Second, one may ask whether analogous packet–level delocalization phenomena hold on nonconvex domains or on Riemannian manifolds with boundary, where boundary Weyl remainders are less rigid.
+Third, the sharpness of the enhanced threshold $k^{(d-3)/d}$ under the double vanishing moment condition, and of the threshold hierarchy in higher dimensions, remains to be investigated.
 Finally, the present comparison principle may serve as a useful input in the study of boundary spectral statistics and correlations, complementing existing microlocal and dynamical approaches.